OpenStudy (anonymous):

find the length of the radius, the circumference, and the area of the circle. use 3.14 for pi. the diameter is 29.7cm (simplify your answer. type an integer or decimal)

OpenStudy (marihelenh):

First, radius is half of the diameter, so do 29.7/2 to get the radius.

OpenStudy (anonymous):

i know c=2xr*3.14 the area=3.14*r^2

OpenStudy (anonymous):

\[\color{green}{1. \quad Radius = \frac{Diameter}{2} \\ 2. \quad Circumference = 2 \times \pi \times Radius \\ 3. \quad Area = \pi \times (Radius)^2}\]

OpenStudy (anonymous):

You need to find Radius to find the other two... Find the radius first of all @marihelenh told you to...

OpenStudy (anonymous):

how to simplify if the radius is 14.85

OpenStudy (marihelenh):

Then the equation for circumference is C=2(pi)r and area equals A=(pi)r^2.

OpenStudy (anonymous):

You can take it as 14.85 itself...

OpenStudy (anonymous):

Do the approximation at the last...

OpenStudy (anonymous):

WOULD IT be cm or cm^2

OpenStudy (marihelenh):

Circumference is cm. Area is cm^2.

OpenStudy (anonymous):

or cm3

OpenStudy (anonymous):

For Circumference, it will be cms only.. And for area, cm^2..

OpenStudy (anonymous):

so the circumference is 93.258

OpenStudy (anonymous):

would the radius be cm or cm2 or cm3

OpenStudy (anonymous):

Radius will be in cms, diameter will be in cms, circum in cms and area will be in cm^2..

OpenStudy (anonymous):

so my answer for the area is 692.44065 how to simplify it the circumference is 93.258 how to simplify that

OpenStudy (marihelenh):

The problem should tell you how it wants you to simplify. If it doesn't say, you probably do not need to worry about simplifying.

OpenStudy (anonymous):

ok it just says simplify (simplify your answer. type an integer or decimal)